OpenStudy (anonymous):

PLEASEEE HELPPP!!! Indicate in standard form the equation of the line passing through the given point and having the given slope. A(5, 5), m = 3

OpenStudy (anonymous):

I got 3x-10=y and it says its wrong

OpenStudy (anonymous):

@sirm3d can you please help?

OpenStudy (goten77):

hmm im not sure whats wrong.. maybe something was typed wrong? or a question error it shud b like this y=mx+b and were given everything but b 5=(5)(3) +b solve for b and get 5 =15 + b b= -10

OpenStudy (anonymous):

but how would that look like in standard form?

OpenStudy (goten77):

oh um

OpenStudy (goten77):

aX + bY = c

OpenStudy (goten77):

y -3x = -10

OpenStudy (anonymous):

it says its wrong..

OpenStudy (goten77):

......what does it say the answer is?

OpenStudy (anonymous):

it doesnt tell me the answer.. it just says its wrong.

OpenStudy (goten77):

where u getting this "telling u wrong" things from?

OpenStudy (anonymous):

It marks it red which means its wrong.

OpenStudy (goten77):

where at? its mostly likely some internet thing that wants it typed some random way

OpenStudy (anonymous):

Its an online course and it has a fill in the blank that looks like this __ - ____

OpenStudy (goten77):

hmm let me think

OpenStudy (goten77):

tbh all i can say is that... the program ur using is the problem. has 2 b a glitch or something wrong w/ how it receives answers.... im 99.9% confident we have the answer

OpenStudy (anonymous):

okay thanks do you think you would know this one ? Indicate in standard form the equation of the line passing through the given point and having the given slope. C(0, 4), m = 0

OpenStudy (goten77):

y=4 since there is no x slope meaing it dont have a x in the equasion so then b is the only thing solved for and b = 4

OpenStudy (anonymous):

in standard form?

OpenStudy (goten77):

it b just y = 4... unless ur program wants to writte out y - 0x = 4 but 0 x = 0

OpenStudy (anonymous):

so y-x=4?

OpenStudy (goten77):

no cause y-x means x has a slope... of 1 and not 0

OpenStudy (anonymous):

ohhh okay